OpenStudy (anonymous):

How many atoms are there in 95.3 g of barium chloride (BaCl2)?

OpenStudy (abb0t):

multiply the given mass by the molar mass. then, proceed to multiply with avogadro's number: \(\sf \color{red}{6.022 \times 10^{23} \frac{atoms}{mol}}\)

OpenStudy (abb0t):

Molar mass = mass of barium \(\times\) \(\sf \color{Red}{2}\)(mass of chlorine) the masses can be found on periodic table under the element.

OpenStudy (abb0t):

You can find their masses on here: http://ptable.com/ just click on any element that you need and the mass comes up.

OpenStudy (anonymous):

so it would be 7.9X10^27

OpenStudy (anonymous):

it was incorrect

OpenStudy (abb0t):

You're close, but I think your power is wrong.

OpenStudy (abb0t):

It should not be 10\(^{27}\)

OpenStudy (anonymous):

i keep getting 27;O

OpenStudy (abb0t):

Try 10\(^{21}\)

OpenStudy (anonymous):

still incorrect:(

OpenStudy (anonymous):

yeah its not right still but thank you for your help!

OpenStudy (abb0t):

2.75 \(\times\) 10\(^{23}\)

OpenStudy (abb0t):

Sorry, I am using an old non-scientific calculator, I think that should be right now that I'm using my phone.